Suppose The Central Bank Sets The Reserve Requirement Ratio At 5%. The Maximum The Central Bank Is Willing To Lend Is 25% Of Required Reserves, Charging The Gross Real Rate Of Return Of 1. We Assume The Gross Real Rate Of Return On Fiat Money Is 1.02 And The Gross Real Return On Capital Is 1.08.

a) What is the gross real return on deposits?

b) If the supply of fiat money is $10,000, what is the quantity of M1 in this economy?
Answer #1

Answer (a) :-

Gross Real return on deposit will be 1.02, which is equal to the Gross Real Rate Of Return On Fiat Money Is 1.02

Reason :

When a Investor deposits money into a bank account, the interest on the deposit is equal to the gross return.

According to the conditions given above ,Fiat Money which usually deposited in the form of Demand Deposits in the banks.

Answer (b) :-

Given ;

Reserve requirement = 5% = 0.05

Lending ratio = 25% of required reserves = 0.25 ×RR

RR = 0.05×10,000 = $500

M1 = $10,000 + 0.25×RR

M1 = $10,000 + 0.25×$500

M1 = $10,000 + $125

M1 = $10,125

Thus, the the quantity of M1 will be $$10,125